# Pion Valence Structure at Intermediate x in the Residual Field Model

### v1

@@ IV.1.1 Soft $q\bar{q}$ Wave Function
For non-relativistic wave function we use two approximations, one Harmonic Oscillator type wave function
@@
and the other the ground state wave function of positronium.
A number of previous works have used the similar wave functions or a form similar to it Brodsky et al. (1981); Lepage and Brodsky (1980); Gunion et al. (1984); Trawiński (2016a); Pasquini and Rodini (2019) Using the above light-front wave function Eq. 36 for the $q\bar{q}$ system’s response function from Eq.29 one obtains:
@@
$=8\pi^{3}\int^{Q^{2}}d^{2}\mathbf{k}_{\perp}\sqrt{{k_{\perp}^{2}\over 4\beta(1-\beta)}}e^{-B_{V}{k_{\perp}^{2}\over 4\beta(1-\beta)}}={16\pi^{4}\sqrt{\pi}A_{V}^{2}\over B_{V}^{3\over 2}}\beta(1-\beta),$ (39)
where in the last step we neglected by the mass of the valence quarks and and assumed $Q^{2}\gg k_{\perp}^{2}$ (consistent with collinear approximation). In calculation of the above integral we set $Q^{2}\rightarrow\infty$.
One can see from the above equation that $q\bar{q}$’s response function peaks at $\beta={1\over 2}$
@@ IV.1.2 Light-Front Wave Function of Valence-Residual System
To calculate the response function of residual system one has to model the the light-front wave function of valence-residual system. Here we use a Harmonic Oscillator (HO) type wavefunction in Eq.(34) obtaining
$\Psi_{12}^{LF}(x_{R},k_{R,\perp})=A_{R}\sqrt{16\pi^{3}}\sqrt{{E_{V}E_{R}\over E_{V}+E_{R}}}e^{-{B_{R}\over 2}k_{R}^{2}},$ (40)
where $E_{V}=\sqrt{m_{V}^{2}+k_{R}^{2}}$, $E_{R}=\sqrt{m_{R}^{2}+k_{R}^{2}}$. The $k_{R}$ is the 3-dimensional momentum defined in the center of mass of the $V-R$ system:
$k^{2}_{VR}=\frac{(s_{VR}-(m_{V}-m_{R})^{2})(s_{VR}-(m_{V}+m_{R})^{2})}{4s_{VR}},$ (41)
@@
Using above expressions for the $V-R$ wave function and relative momentum $k_{R}$, for the response function of the $VR$ system one obtains:
@@
Note that in the above equation masses of the valence cluster $m_{V}$, and and residual state $m_{R}$, are virtual and characterize the transition into non-stationary virtual state of a valence cluster and the residual system, thus they do not necessarily need to correspond to the pion mass. Additionally there is no condition for collinear approximation so the response function may exhibit some $Q^{2}$ dependence.
@@ IV.2 Modeled Soft Pion PDF
Inserting Eq.(39) and Eq.44) into Eq.(29) for the modeled parton distribution of pion one obtains:
$f_{V}(x_{B},Q^{2})={(\pi)^{3\over 2}A_{V}^{2}A_{R}^{2}\over B_{V}^{3\over 2}}\int_{0}^{1-x_{B}}\frac{dx_{R}}{x_{R}}{1\over(1-x_{R})^{2}}\ \left[\int^{Q^{2}}d^{2}\mathbf{k}_{R,\perp}{E_{V}E_{R}\over(E_{V}+E_{R})}e^{-B_{R}k_{R}^{2}}\right]$
V Fitting Results
V.1 Parameters of Fitting
We use recent results from JAM18 collaborationBarry et al. (2018b, 2021b) on pion PDF to fit Eq.(45). Our main focus in the fitting procedure is to describe the position and the hight of the x weighted parton distribution in accordance with our model assumption that these features of valence PDFs are characterized by the soft partonic dynamics. As it follows from Eq.(45) the masses of residual, $m_{R}$, and valence, $m_{V}$, systems, as well as the width of the $V-R$ distribution define the position and the height of the peak.
We choose starting $Q^{2}=Q_{0}^{2}=1.69$ GeV2, where $Q_{0}$ is chosen in JAM18 parameterization to be the $c$-quark mass. The extensive fitting analysis indicates that the position of the peak can be reproduced only if the virtual mass of the residual system is close to zero. This is very different from our analysis of nucleon PDF’s Leon and Sargsian (2020); Leon et al. (2020) in which we found that the residual mass for the best fit is close to the pion mass, $m_{R}^{N}\approx 0.14$, indicating the validity of meson cloud picture of valence parton distribution in the nucleon.
The best fit of the JAM18 pion PDF is presented in Fig.2, for the parameters of $m_{R}=0$, $m_{V}=0.95$ GeV and $B_{V}=1.2$GeV-2. Note that the qualitative agreement of the data requires $m_{V}\sim Q_{0}$ and $m_{R}\sim 0$. As the figure shows, the soft mechanism describes the empirical PDF reasonably well in the range of $0.850.1<x<$ 0.85, leaving very little room for the hard mechanism for the generation of PDF at large x.
@@ V.2 Interpretation of the Results
The parameters of the model that provide the best fit to the empirical PDF allows to accertain the mechanism dominating parton dynamics at the range where the peak of x-weighted PDF distribution is observed. For this we first analyze the response function of the residual system presented in Eq.(44). As Fig.3 shows the response function has a wide distribution at $x_{R}$ slightly peaking at $x_{R}\approx 0.5$.
@@
This indicates that the valence $q\bar{q}$ system carries a total momentum fraction $1-x_{R}\approx 0.5$. Since $x$-Bjorken corresponds to the Light-Front momentum fraction of interacting quark, then one concludes that the observed distribution indicates the dominance of the Feynman mechanismFeynman (1972) according to which the interacting quark carries the most of the Light-front momentum fraction with remaining momenta distributed among wee partons.
Another observation, as it was noted above, the result of $m_{R}\approx 0$, this indicates that the residual system does not exhibit any structure (as it was case of nucleon PDFs). On the other hand the large $m_{V}\approx 0.95$ GeV indicate the large virtuality of interacting quark comparable with the $Q_{0}$ of the probe. This together with the relatively small $B_{R}$ indicates that while interacting valence quark is highly virtual, the residual system is spread over the bulk size of the pion.
This picture is consistent with the phenomenology that nucleon interactions in a nucleus proceed through the pion exchange in which no quark-interchanges are involved.
Finally, the surprising result of the fit is that the considered model leaves little room for hard mechanism of high x PDF generated though the hard gluon exchange between valence $q\bar{q}$ system.
@@ VI Conclusion and Outlook
We have calculated pion valence parton distribution within residual field model, which was previously applied to the calculation of Nucleon PDFs. The models assume no hard interaction between valence quarks and it separates $q\bar{q}$ valence cluster from the residual system. Both, $q\bar{q}$ and $V-R$ systems are described by nonperturbative Light-Front wave functions that satisfies relativistic normalization. These wavefunctions allow to introduce the residual response function as well as the response function of the $q\bar{q}$ valence cluster. The pion PDF represents the convolution of these response functions. We analyzed the parameters of the model by fitting the position and the height of the x-weighted PDF using empirical JAM18 distribution.
The fitting indicates that one can describe the above attributes of pion PDF only if the mass of the residual system is close to zero. However we obtain large virtual mass for the $q\bar{q}$ cluster comparable to the $Q_{0}$ of the probe. Analysis of the response function of the residual system indicates that the dominating mechanism is similar to the Feynman mechanism in which interacting parton carries almost all the momentum of the valence quark system, with the remaining light-front momentum distributed to the non-interacting wee partons.
The fit allows to describe the empirical PDF reasonably well in the range of $0.1\leq x\leq 0.85$ leaving little room for the hard mechanism of generating PDFs at large x. The model practically resolves the observed $(1-x)^{\beta}$, $\beta\approx 1$ scaling relation of the pion PDFs at large x.
With the parameters of $q\bar{q}$ and residual system Light-Front wave functions fixed one will be able to make numerical predictions for different observables such as generalized parton distribution functions (GPDs) or transverse momentum dependent parton distribution functions (TMDs) in pion that will allow for additional verification of the model.
@@ Acknowledgement
This work is supported by United States Department of Energy grant under contract DE-FG02-01ER41172. AS acknowledges support through a Bridge to Doctorate NSF fellowship at FIU.
Appendix A Mandelstam s with many particles
For $n$ particles each with 4-momentum
$k^{\mu}_{i}=\left(k^{+}_{i},\frac{k_{\perp,i}^{2}+m^{2}_{i}}{k_{i}^{+}},\mathbf{k}^{\perp}_{i}\right)$ (46)
and total 4-momentum $k^{\mu}_{T}=\sum_{i=1}^{n}k^{\mu}_{i}=(x_{T}P^{+},k^{-}_{T},\mathbf{k}^{\perp}_{T})$, then the generalized Mandelstam $s_{n}$ is:
$s_{n}=\Bigg{(}\sum_{i}^{n}k^{\mu}_{i}\Bigg{)}^{2}=\sum_{i}^{n}\frac{\tilde{k}_{i,\perp}^{2}+m_{i}^{2}}{\beta_{i}}$ (47)
where:
$\beta_{i}=\frac{x_{i}}{x_{T}}$ (48)
$\mathbf{\tilde{k}}_{i,\perp}=\mathbf{k}_{i,\perp}-\beta_{i}\mathbf{k}_{T,\perp}$ (49)
The is seen by:
$s_{n}$ $=\Bigg{(}\sum_{i}^{n}k^{\mu}_{i}\Bigg{)}^{2}=\Bigg{(}\sum_{i}^{n}k^{+}_{i}\Bigg{)}\Bigg{(}\sum_{i}^{n}k^{-}_{i}\Bigg{)}-\Bigg{(}\sum_{i}^{n}\mathbf{k}_{i,\perp}\Bigg{)}^{2}$ (50)
$=(x_{T}P^{+})\Bigg{(}\sum_{i}^{n}\frac{{k}_{i,\perp}^{2}+m_{i}^{2}}{x_{i}P^{+}}\Bigg{)}-k^{2}_{T,\perp}$ (51)
$=\sum_{i}\frac{k^{2}_{i,\perp}+m^{2}_{i}}{\frac{x_{i}}{x_{T}}}-k^{2}_{T,\perp}$ (52)
Using the definition of $\beta_{i}$ and $\tilde{\mathbf{k}}_{i,\perp}$:
$s_{n}$ $=\sum_{i}\frac{\tilde{k}^{2}_{i,\perp}+2\beta_{i}\mathbf{k}_{T,\perp}\cdot\tilde{\mathbf{k}}_{i,\perp}+\beta_{i}^{2}k_{T,\perp}+m^{2}_{i}}{\beta_{i}}-k^{2}_{T,\perp}$ (53)
$=\sum_{i}\frac{\tilde{k}^{2}_{i,\perp}+2\beta_{i}\mathbf{k}_{T,\perp}\cdot\tilde{\mathbf{k}}_{i,\perp}+\beta_{i}^{2}k_{T,\perp}+m^{2}_{i}}{\beta_{i}}-k^{2}_{T,\perp}$ (54)
$=\sum_{i}\Bigg{(}\frac{\tilde{k}^{2}_{i,\perp}+m^{2}_{i}}{\beta_{i}}+2k_{T,\perp}\cdot\tilde{k}_{i,\perp}+\beta_{i}k^{2}_{T,\perp}\Bigg{)}-k^{2}_{T,\perp}$ (55)
Using the fact that $\sum_{i}\beta_{i}=1$ and $\sum_{i}\tilde{\mathbf{k}}_{i,\perp}=\sum_{i}({\mathbf{k}}_{i,\perp}-\beta_{i}\mathbf{k}_{T})=\mathbf{0}$,
$s_{n}=\sum_{i}\Bigg{(}\frac{\tilde{k}^{2}_{i,\perp}+m^{2}_{i}}{\beta_{i}}+k^{2}_{T,\perp}-k^{2}_{T,\perp}\Bigg{)}$ (56)
Then,
$s_{n}=\sum_{i}\frac{\tilde{k}^{2}_{i,\perp}+m^{2}_{i}}{\beta_{i}}$ (57)
Appendix B Light Front Energy Denominators
The LF energy denominator, $D$, for an intermediate state is given by:
$D$ $=\sum(k^{-}_{initial}-k^{-}_{intermediate})$ (58)
where $+$ and transverse momenta are conserved and we treat each particle on-mass shell, $k^{-}=\frac{m^{2}+k^{2}_{\perp}}{k^{+}}$. The denominator for nucleon splitting to the residual and valence parts is $D_{1}$:
$D_{1}$ $=p_{N}^{-}-k_{R}^{-}-k_{V}^{-}$
$=\frac{M^{2}_{N}}{p_{N}^{+}}-\frac{k_{R,\perp}+m^{2}_{R}}{k^{+}_{R}}-\frac{k_{V,\perp}+m^{2}_{V}}{k^{+}_{V}}$
$=\frac{1}{p^{+}_{N}}\left(M_{N}^{2}-\frac{k_{R,\perp}+m^{2}_{R}}{x_{R}}-\frac{k_{V,\perp}+m^{2}_{V}}{x_{V}}\right)$ (59)
Where $x_{R}=\frac{k^{+}_{R}}{p_{N}^{+}}$ and $x_{V}=\frac{k^{+}_{V}}{p_{N}^{+}}.$ The denominator for valence splitting to the 2 valence quarks of the Pion is $D_{2}$:
$D_{2}$ $=k^{-}_{V}+k^{-}_{R}-(k^{-}_{1i}+k^{-}_{2i}+k^{-}_{R})$
$=\frac{m^{2}_{V}+k_{V,\perp}^{2}}{k^{+}_{V}}-\frac{m^{2}_{1}+k_{1i,\perp}^{2}}{k^{+}_{1i}}-\frac{m^{2}_{2}+k_{2i,\perp}^{2}}{k^{+}_{2i}}$
$=\frac{1}{k^{+}_{V}}\left(m_{V}^{2}+k_{V,\perp}^{2}-\frac{k^{2}_{1i,\perp}+m^{2}_{1}}{\beta_{1i}}-\frac{k^{2}_{2i,\perp}+m^{2}_{2}}{\beta_{2i}}\right)$ (60)
Where $\beta_{1i}=\frac{k^{+}_{1i}}{k^{+}_{V}}$ and, $\beta_{2i}=\frac{k^{+}_{2i}}{k^{+}_{V}}$The denominator of the quark - quark interaction is $D_{3}$:
$D_{3}$ $=k^{-}_{1i}+k^{-}_{2i}-(k^{-}_{1m}+k^{-}_{2})$
$=\frac{1}{p_{N}^{+}}\Bigg{(}\frac{m^{2}_{1}+k_{1i,\perp}^{2}}{x_{1i}}+\frac{m^{2}_{2}+k_{2i,\perp}^{2}}{x_{2i}}$
$-\frac{m^{2}_{1}+k_{1m,\perp}^{2}}{x_{1m}}-\frac{m^{2}_{2}+k_{2,\perp}^{2}}{x_{2}}\Bigg{)}$
$=\frac{1}{p_{N}^{+}}\tilde{D_{3}}$ (61)

### v2

@@ IV.1.1 Soft $q\bar{q}$ Wave Function
We normalize the $q\bar{q}$ wave function per one quark (or anitquark) in the cluster since in the calculation of the PDF we sum by individual contributions of valence quarks. We also take into account that the normalization is less than unity since no hard component is included in the cluster wave function which can be generated through the hard gluon exchangesLepage and Brodsky (1980). With this, the normalizetion is fixed as:
$\int\mid\Psi^{LF}{q\bar{q}}(k_{12})\mid^{2}{d\beta\over\beta(1-\beta)}{d^{2}k_{12,\perp}\over 16\pi^{3}}=N_{q\bar{q}}=1-h.$ (38)
Note that in addition to the hard component the parameter $h$ also accounts for the fact that at $x<0.1$, the Regge dynamics may play a role in PDF which is not included in current calculations.
For non-relativistic wave function we use a Harmonic Oscillator type wave function with free parameters, $A_{V}$ and $B_{V}$:
@@
A number of previous works have used the similar wave functions or a form similar to it Brodsky et al. (1981); Lepage and Brodsky (1980); Gunion et al. (1984); Trawiński (2016a); Pasquini and Rodini (2019). Using the above light-front wave function, Eq.(36), for the $q\bar{q}$ system’s response function defined in Eq.(29), one obtains:
@@
$=8\pi^{3}\int^{Q^{2}}d^{2}\mathbf{k}_{\perp}\sqrt{{k_{\perp}^{2}\over 4\beta(1-\beta)}}e^{-B_{V}{k_{\perp}^{2}\over 4\beta(1-\beta)}}={16\pi^{4}\sqrt{\pi}A_{V}^{2}\over B_{V}^{3\over 2}}\beta(1-\beta)=16\pi^{3}N_{q\bar{q}}\beta(1-\beta).$
Here in the last step we neglected by the masses of the valence quarks and assumed $Q^{2}\gg k_{\perp}^{2}$ (consistent with collinear approximation). This allowed to calculate the integral analytically, in which we set the upper limit of the integration to $\infty$. It is interesting that in the massless limit because of the relation of Eq.(37) the integral in the r.h.s part of the equation has an analytic form of the normalization of the NR wave function in the center of mass of $q\bar{q}$ system. As a result the response function depends on the overall normalization factor $N_{q\bar{q}}={\pi^{3\over 2}A_{V}^{2}\over B_{V}^{3\over 2}}$ while momentum distribution, $\beta(1-\beta)$ is similar to the one obtained from perturbative wave function.
It is worth mentioning that the calculated in this way $q\bar{q}$’s response function peaks at $\beta={1\over 2}$ which is consistent with relativistic wave function of massless $q\bar{q}$ system.
@@ IV.1.2 Light-Front Wave Function of Valence-Residual System
To calculate the response function of residual system one has to model the light-front wave function of valence-residual system. Here we use a Harmonic Oscillator (HO) type wave function in Eq.(34) obtaining
$\Psi_{R}^{LF}(x_{R},k_{R,\perp})=A_{R}\sqrt{16\pi^{3}}\sqrt{{E_{V}E_{R}\over E_{V}+E_{R}}}e^{-{B_{R}\over 2}k_{R}^{2}},$ (41)
where $E_{V}=\sqrt{m_{V}^{2}+k_{R}^{2}}$, $E_{R}=\sqrt{m_{R}^{2}+k_{R}^{2}}$. The $k_{R}$ is the 3-dimensional momentum defined in the center of mass of the VRV-R system:
$k^{2}_{R}=\frac{(s_{VR}-(m_{V}-m_{R})^{2})(s_{VR}-(m_{V}+m_{R})^{2})}{4s_{VR}},$ (42)
@@
The above defined wave function is normalized as follows:
$\int\mid\Psi^{LF}_{R}(k_{12})\mid^{2}{dy\over x_{R}(1-x_{R})}{d^{2}k_{R,\perp}\over 16\pi^{3}}=N_{R},$ (45)
where the normalization, $N_{R}$ one can interpret as the number of constituents in the residual hadronic system.
Using the above expressions for the $VR$ wave function and the relative momentum $k_{R}$ (Eq.(42)), for the response function of the $VR$ system, one obtains the following:
@@
Note that in the above equation masses of the valence cluster $m_{V}$, and residual state $m_{R}$, are virtual and characterize the transition into non-stationary virtual state of a valence cluster and the residual system, thus they do not need to correspond to the pion mass.
@@ IV.2 Modeled Soft Pion PDF
Inserting Eq.(40) and Eq.(46) into Eq.(29) and using normalization codition of Eq.(38, for the modeled parton distribution of pion one obtains:
$f_{V}(x_{B},Q^{2})={N_{q\bar{q}}A_{R}^{2}}\int_{0}^{1-x_{B}}\frac{dx_{R}}{x_{R}}{1\over(1-x_{R})^{2}}\ \left[\int^{Q^{2}}d^{2}\mathbf{k}_{R,\perp}{E_{V}E_{R}\over(E_{V}+E_{R})}e^{-B_{R}k_{R}^{2}}\right].$
The obtained expression indicates that, the $\beta(1-\beta)$- relative light-cone momentum fraction distribution of valence quarks in the cluster (Eq.(40)) results in an only $x_{R}$ dependent integrand in Eq.(47) convoluted with the response function of the residual system (Eq(46)). In the current model this follows from the neglecting the valence quark masses. As Eq(47) shows the shape of the PDF depends on the slope factor of the cluster-residual system’s momentum distribution, $B_{R}$ as well as the respective virtual masses of the cluster $m_{V}$ and the residual system $m_{R}$.
V Numerical Results
V.1 Parameters of the model
We use recent results from JAM collaborationBarry et al. (2018b, 2021b) on pion PDF to fit Eq.(47). Our main focus in the fitting procedure is to describe the position and the height of the x weighted parton distribution in accordance with our model assumption that these features of valence PDFs are characterized by the soft partonic dynamics. As follows from Eq.(47) the masses of residual, $m_{R}$, and valence, $m_{V}$, systems, as well as the width of the $VR$ distribution, $B_{R}$ define the shape of the PDF that produces the peaking distribution of the PDF weighted by $x$.
We choose the starting $Q^{2}=Q_{0}^{2}=1.69$ GeV2, where $Q_{0}$ is chosen in JAM parameterization to be the $c$-quark mass. The fitting analysis indicates that the position of the peak can be reproduced only if the virtual mass of the residual system is close to zero, $m_{R}\approx 0$. This is very different from our analysis of nucleon PDF’s in residual field modelLeon and Sargsian (2020); Leon et al. (2020) in which we found that the residual mass for the best fit is close to the pion mass, $m_{R}^{N}\approx 0.14$, indicating the validity of meson cloud picture of valence parton distribution in the nucleon.
The best fit of the JAM pion PDF is presented in Fig.2, for the parameters of $m_{R}=0$, $m_{V}=0.95$ GeV and $B_{V}=1.2$GeV-2. Note that the qualitative agreement of the data requires $m_{V}\sim Q_{0}$ and $m_{R}\sim 0$. As the figure shows, the soft mechanism describes the empirical PDF reasonably well in the range of $0.1<x<0.85$, leaving very little room for the hard mechanism in the generation of PDF at large x.
@@ V.2 Interpretation of the Results
The parameters of the model that provide the best fit to the empirical PDF allow to ascertain the mechanism dominating parton dynamics at the x-range where the peak of x-weighted PDF distribution is observed. For this, we first analyze the response function of the residual system presented in Eq.(46). As Fig.3 shows the response function at starting $Q_{0}$, has a wide distribution in $x_{R}$, slightly peaking at $x_{R}\approx 0.5$. The distribution saturates at higher $Q^{2}$ with the peaking at $x_{R}=0.5$ becoming more pronounced. This indicates that the valence $q\bar{q}$ system as a whole carries a total momentum fraction $x_{V}=1-x_{R}\approx 0.5$.
@@
Since $x$-Bjorken corresponds to the light-front momentum fraction of the interacting quark with the relation $x\approx\beta(1-x_{R})$, where $\beta$ is the LF momentum fraction of the $q\bar{q}$ cluster carried by the interacting quark, and since the broad distribution of $x_{R}$, one observes that for $x\gtrsim 0.5$, relative momentum fraction, $\beta\sim 1$. This indicates that the observed distribution favors the dominance of the Feynman mechanismFeynman (1972) according to which the interacting quark carries the most of the light-front momentum fraction of the cluster, with remaining LF momenta distributed among the wee partons in the residual system.
Another observation (as noted above) is the result of $m_{R}\approx 0$. This indicates that the residual system does not exhibit any structure (as it was the case for valence PDFs of nucleons Leon and Sargsian (2022)). On the other hand, the large value of $m_{V}\approx 0.95$ GeV indicates the large virtuality of the interacting quark comparable to $Q_{0}$ of the probe. This together with the relatively large $B_{R}$ indicates that while interacting valence quark is highly virtual, the residual system is spread over the bulk size of the pion. This picture is consistent with the phenomenology that nucleon-nucleon interactions at large distances ($\geq 1.5fm$) proceeds through the pion exchange in which no quark-interchanges are involved.
The overall constant in Eq.(47), is evaluated at $\sqrt{N_{q\bar{q}}}A_{R}=\sqrt{1-h}A_{R}\approx 0.86$. Assuming that overall Regge and high momentum mechanism contribute to $h\approx 0.2$, one obtains for the normalization of the residual LF wave function in Eq.(45) $N_{R}\approx 4$. This will predict, for example, that integrated multiplicity for pion in the target fragmention region to be above $2$.
Finally, the surprising result of the fit is that the considered model leaves little room for hard mechanism of high x PDF generated through the hard gluon exchange between valence $q\bar{q}$ system. The current result favors a pion structure dominated by soft mechanism of valence quark interaction. Moreover, it is interesting that $f_{V}(x)\mid_{x\to 1}\sim(1-x)$ consistent with the $f_{V}(x)\sim 1-x$ distribution at the intermediate region in which case $xf_{V}(x)\sim x(1-x)$ peaks at $x_{p}\sim 0.5$ in agreement with the data. The later reinforces that observation that high $x$ structure of pion PDF has large soft component that defines also the peaking properties of $xf_{V}(x)$ distribution.
@@ VI Conclusion and Outlook
We have calculated pion valence parton distribution within residual field model, which was previously applied to the calculation of nucleon PDFs. The model assumes no hard interaction between valence quarks and it separates $q\bar{q}$ valence cluster from the residual system. Both, $q\bar{q}$ and $VR$ systems are described by nonperturbative light-front wave functions that satisfies relativistic normalization. These wavefunctions allow to introduce the residual response function as well as the response function of the $q\bar{q}$ valence cluster. The pion PDF represents the convolution of these response functions. We analyzed the parameters of the model by fitting the position and the height of the x-weighted PDF using empirical JAM18 distribution.
The fitting indicates that one can describe the above attributes of pion PDF only if the mass of the residual system is close to zero. However, we obtain large virtual mass for the $q\bar{q}$ cluster comparable to the $Q_{0}$ of the probe. Analysis of the response function of the residual system indicates that the dominant mechanism is similar to the Feynman mechanism in which interacting parton carries almost all the momentum of the valence quark system, with the remaining light-front momentum distributed to the non-interacting wee partons.
The fit allows to describe the empirical PDF reasonably well in the range of $0.1\leq x\leq 0.85$ leaving little room for the hard mechanism of generating PDFs at large x. The model practically resolves the observed $(1-x)^{\beta}$, $\beta\approx 1$ scaling relation of the pion PDFs at large x. The large soft part contribution at high x, indicates that the introduction of the hard gluon exchange mechanism in the generation of high x PDF will require the consideration of the interference between soft and hard amplitudes that can introduce the exponent, $\beta<2$ in $(1-x)^{\beta}$ distribution at very large x.
With the parameters of light-Front wave function of residual system fixed, one will be able to make numerical predictions for different observables such as electromagnetic form-factors, generalized parton distribution functions (GPDs) or transverse momentum dependent parton distribution functions (TMDs) in pion that will allow an additional verification of the model. Additionally, the modeled residual state wave function allows to evaluate the fragmentation of pion in the backward region that can be probed in the EIC collider kinematics.
@@ Acknowledgement
Authors are thankful to Alan Sosa for useful discussions. This work is supported by United States Department of Energy grant under contract DE-FG02-01ER41172.